Evaluating a regression model: A regression was run to determine if there is a relationship between hours of TV watched per day (x) and number of situps a person can do (y). The results of the regression were: y = − 0.693 ⋅ x + 28.802 , with an R-squared value of 0.571536. Assume the model indicates a significant relationship between hours of TV watched and the number of situps a person can do. Use the model to predict the number of situps a person who watches 6 hours of TV can do (to one decimal place).
Answer #1

Solution:

The given regression model is

Y= 28.802 - 0.693*X

Now we have to predict Y for x=6, we get

Y= 28.802- 0.693*X

Y= 28.802 - 0.693*6

Y= 24.6

• The predicted number of situps a person who watches 6 hours of tv is 24.6